OpenStudy (anonymous):

solve. X+10/ x squared -2 = 4/x

OpenStudy (anonymous):

\[\frac{x+10}{x^2}-2=\frac{4}{x}\]?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

x squared -2 is the denominator

OpenStudy (anonymous):

or \[\frac{x+10}{x^2-2}=\frac{4}{x}\]

OpenStudy (anonymous):

cross multiply and solve the resulting quadratic equation. first step is \[x(x+10)=4(x^2-2)\]

OpenStudy (anonymous):

yes this one

OpenStudy (anonymous):

\[x^2+10x=4x^2-8\] \[3x^2-10x-8=0\] and if you are lucky it factors and if not use the formula

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[(x-4)(3x+2)=0\] etc